# Robust Semantic Communications for Speech Transmission

### v2

@@ III-B Ross-S2T
As aforementioned, the deep semantic compensator is responsible for estimating the damaged semantic information in $\overline{\boldsymbol{S}}$, extracting the deep semantic features with the least dissimilarity to $\boldsymbol{F}$, and returning the semantic impairment probe matrix to record the positional information of corrupted deep semantic features. In the second training stage, we propose a dual-compensator mechanism, including the GAN-enabled deep semantic compensator at the transmitter and the probe-aided compensator at the receiver. Particularly, the trained deep semantic encoder is leveraged to obtain $\boldsymbol{F}$ as the real data for the discriminator. The generator is developed to process the corrupted speech, $\overline{\boldsymbol{S}}$, and generate the fake data $\widetilde{\boldsymbol{F}}$ to fool the discriminator by adopting the 1D CNN module followed by the latent space, $\boldsymbol{Z}$, and the transformer module. Note that the intermediate semantic representation, $\widetilde{\boldsymbol{I}}$, are attained as the output of the 1D CNN module. The discriminator distinguishes whether the input data is real or fake, which incorporates the 2D CNN module prior to latent space $\boldsymbol{Z}$ followed by the dense layer and the sigmoid layer. Denote the trainable NN parameters of the discriminator and the generator as ${\boldsymbol{\gamma}}_{\mathrm{D}}$ and ${\boldsymbol{\gamma}}_{\mathrm{G}}$, respectively, i.e., $\boldsymbol{\gamma}=({\boldsymbol{\gamma}}_{\mathrm{D}},\;{\boldsymbol{\gamma}}_{\mathrm{G}})$. Then, the loss function adopted for training the discrimination can be expressed as
${\mathcal{L}}_{\mathrm{D}}(\boldsymbol{S},\;\overline{\boldsymbol{S}};\;{\boldsymbol{\gamma}}_{\mathrm{D}})=\frac{1}{2}\left({\mathfrak{T}}_{\mathrm{D}}\left({\mathfrak{T}}_{\mathrm{S}}\left(\boldsymbol{S}\right)\right)-1\right)^{2}+\frac{1}{2}\left({\mathfrak{T}}_{\mathrm{D}}({\mathfrak{T}}_{\mathrm{G}}(\overline{\boldsymbol{S}}))\right)^{2},$ (8)
where ${\mathfrak{T}}_{\mathrm{D}}(\cdot)$ and ${\mathfrak{T}}_{\mathrm{G}}(\cdot)$ are the discriminator and the generator, respectively.
Furthermore, the ${\boldsymbol{\gamma}}_{\mathrm{G}}$ can be updated through the loss function as follows,
$\begin{split}{\mathcal{L}}_{\mathrm{G}}(\boldsymbol{S},\;\overline{\boldsymbol{S}};\;{\boldsymbol{\gamma}}_{\mathrm{G}})&=\frac{1}{2}\xi{\mathcal{L}}_{\mathrm{MSE}}+\frac{1}{2}{\mathcal{L}}_{\mathrm{ADV}}\\ &=\frac{1}{2}\xi\left(\boldsymbol{F}-\widetilde{\boldsymbol{F}}\right)^{2}+\frac{1}{2}\left({\mathfrak{T}}_{\mathrm{D}}({\mathfrak{T}}_{\mathrm{G}}(\overline{\boldsymbol{S}}))-1\right)^{2},\end{split}$ (9)
@@
As a result, the trained generator can create realistic data to deceive the discriminator and produce compensated deep semantic features, $\widetilde{\boldsymbol{F}}$, that closely resemble $\boldsymbol{F}$. The details for training the deep semantic compensator are presented in Algorithm 2.
Furthermore, a probe network is established at the transmission by taking $\widetilde{\boldsymbol{I}}$ as the input and providing the semantic impairment probe, $\boldsymbol{C}$, written as
$\boldsymbol{C}=\mathfrak{T}_{\mathrm{PN}}(\widetilde{\boldsymbol{I}})\;\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\delta},$ (10)
where $\mathfrak{T}_{\mathrm{PN}}(\cdot)$ is the probe network and $\boldsymbol{\delta}$ is NN parameters.
Algorithm 1 Training algorithm of DeepSC-S2T.
Initialization: Initialize trainable parameters $\boldsymbol{\theta}$.
1:Input: Clear speech input $\boldsymbol{S}$ and accurate target text $\widetilde{\boldsymbol{T}}$ from tainset $\boldsymbol{\mathfrak{S}}$, fading channel $\boldsymbol{H}$, Gaussian noise $\boldsymbol{N}$.
2:while loss ${\mathcal{L}}_{\mathrm{LSR-CE}}(\boldsymbol{\theta})$ is not converged do
3:     ${\mathfrak{T}}_{\mathrm{C}}({\mathfrak{T}}_{\mathrm{S}}(\boldsymbol{S}))\rightarrow\boldsymbol{X}$.
4:     Transmit $\boldsymbol{X}$ over $\boldsymbol{H}$ and receive $\boldsymbol{Y}$ via (2).
5:     $\mathfrak{T}_{\mathrm{S}}^{-1}(\mathfrak{T}_{\mathrm{C}}^{-1}\left(\boldsymbol{Y}\right))\rightarrow\widehat{\widetilde{\boldsymbol{T}}}$.
6:     Compute ${\mathcal{L}}_{\mathrm{LSR-CE}}(\boldsymbol{\theta})$ to update $\boldsymbol{\theta}$.
7:end while
8:Output: Trained ${\mathfrak{T}}_{\mathrm{S}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{C}}(\cdot)$, $\mathfrak{T}_{\mathrm{C}}^{-1}(\cdot)$, and $\mathfrak{T}_{\mathrm{S}}^{-1}(\cdot)$.
The loss function for training the $\mathfrak{T}_{\mathrm{PN}}(\cdot)$ is denoted as
${\mathcal{L}}_{\mathrm{PN}}(\boldsymbol{I},\;\widetilde{\boldsymbol{I}},\;\boldsymbol{C};\;\boldsymbol{\delta})=\sum_{l^{{}^{\prime}}=1}^{L^{{}^{\prime}}}\left(i_{l^{{}^{\prime}}}-c_{l^{{}^{\prime}}}{\widetilde{i}}_{l^{{}^{\prime}}}\right)^{2},$ (11)
where $\boldsymbol{I}$ is the intermediate semantic representation extracted from the clear speech.
To further enhance the fidelity of the received deep semantic features, $\widehat{\widetilde{\boldsymbol{F}}}$, the learned semantic impairment probe, $\boldsymbol{C}$, is utilized to identify the corrupted deep semantic feature in $\widehat{\widetilde{\boldsymbol{F}}}$ and the CNN-enabled probe-aided compensator commits to reducing semantic errors between the identified corrupted features and the corresponding accurate features. The compensated deep semantic features, $\widehat{\boldsymbol{F}}$, can be expressed as
$\widehat{\boldsymbol{F}}=\mathfrak{T}_{\mathrm{PC}}(\widehat{\widetilde{\boldsymbol{F}}},\;\boldsymbol{C})\;\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\zeta},$ (12)
where $\mathfrak{T}_{\mathrm{PC}}(\cdot)$ indicates the probe-aided compensator and $\boldsymbol{\zeta}$ is its NN parameters.
Besides, the $\boldsymbol{\zeta}$ can be updated by
@@
where $l^{\backprime}$ indicates the position corresponding to the semantic impairment probe with the value of one.
Algorithm 2 Training algorithm of GAN-enabled deep semantic compensator with corrupted speech input.
Initialization: Initialize trainable parameters ${\boldsymbol{\gamma}}_{\mathrm{D}}$ and ${\boldsymbol{\gamma}}_{\mathrm{G}}$.
1:Input: Clear speech input $\boldsymbol{S}$ and corrupted speech input $\overline{\boldsymbol{S}}$ from tainset $\boldsymbol{\mathfrak{S}}$.
2:Obtain the trained ${\mathfrak{T}}_{\mathrm{S}}(\cdot)$ from Algorithm 1.
3:while losses ${\mathcal{L}}_{\mathrm{G}}({\boldsymbol{\gamma}}_{\mathrm{D}})$ and ${\mathcal{L}}_{\mathrm{G}}({\boldsymbol{\gamma}}_{\mathrm{G}})$ are not converged do
4:     Generate real data via ${\mathfrak{T}}_{\mathrm{S}}(\boldsymbol{S})\rightarrow\boldsymbol{F}$.
5:     Generate fake data via ${\mathfrak{T}}_{\mathrm{G}}(\overline{\boldsymbol{S}})\rightarrow\widetilde{\boldsymbol{F}}$.
6:     Distinguish real data via ${\mathfrak{T}}_{\mathrm{D}}(\boldsymbol{F})$.
7:     Distinguish fake data via ${\mathfrak{T}}_{\mathrm{D}}(\widetilde{\boldsymbol{F}})$.
8:     Compute ${\mathcal{L}}_{\mathrm{D}}({\boldsymbol{\gamma}}_{\mathrm{D}})$ and update ${\boldsymbol{\gamma}}_{\mathrm{D}}$.
9:     for each update of generator ${\mathfrak{T}}_{\mathrm{G}}(\cdot)$ do
10:         Generate real data via ${\mathfrak{T}}_{\mathrm{S}}(\boldsymbol{S})\rightarrow\boldsymbol{F}$.
11:         Generate fake data via ${\mathfrak{T}}_{\mathrm{G}}(\overline{\boldsymbol{S}})\rightarrow\widetilde{\boldsymbol{F}}$.
12:         Distinguish fake data via ${\mathfrak{T}}_{\mathrm{D}}(\widetilde{\boldsymbol{F}})$.
13:         Compute ${\mathcal{L}}_{\mathrm{G}}({\boldsymbol{\gamma}}_{\mathrm{G}})$ to update ${\boldsymbol{\gamma}}_{\mathrm{G}}$.
14:     end for
15:end while
16:Output: Trained ${\mathfrak{T}}_{\mathrm{D}}(\cdot)$ and ${\mathfrak{T}}_{\mathrm{G}}(\cdot)$.
Algorithm 3 Training algorithm of probe network and probe-aided compensator.
Initialization: Initialize trainable parameters $\boldsymbol{\delta}$ and $\boldsymbol{\zeta}$.
1:Input: Corrupted speech input $\overline{\boldsymbol{S}}$ from tainset $\boldsymbol{\mathfrak{S}}$.
2:Obtain the trained ${\mathfrak{T}}_{\mathrm{C}}(\cdot)$, $\mathfrak{T}_{\mathrm{C}}^{-1}(\cdot)$, and $\mathfrak{T}_{\mathrm{S}}^{-1}(\cdot)$ from Algorithm 1, the trained ${\mathfrak{T}}_{\mathrm{G}}(\cdot)$ from Algorithm 2.
3:while loss ${\mathcal{L}}_{\mathrm{PN}}(\boldsymbol{\delta})$ and loss ${\mathcal{L}}_{\mathrm{PC}}(\boldsymbol{\zeta})$ are not converged do
4:     Attain intermedia semantic representation $\widetilde{\boldsymbol{I}}$.
5:     Obtain semantic impairment probe via ${\mathfrak{T}}_{\mathrm{PN}}(\widetilde{\boldsymbol{I}})\rightarrow\boldsymbol{C}$.
6:     Generate deep semantic features via ${\mathfrak{T}}_{\mathrm{G}}(\overline{\boldsymbol{S}})\rightarrow\widetilde{\boldsymbol{F}}$.
7:     ${\mathfrak{T}}_{\mathrm{C}}(\widetilde{\boldsymbol{F}})\rightarrow\boldsymbol{X}$.
8:     Transmit $\boldsymbol{X}$ over $\boldsymbol{H}$ and receive $\boldsymbol{Y}$ via (2).
9:     $\mathfrak{T}_{\mathrm{C}}^{-1}\left(\boldsymbol{Y}\right)\rightarrow\widehat{\widetilde{\boldsymbol{F}}}$.
10:     Compensate $\widehat{\widetilde{\boldsymbol{F}}}$ via $\mathfrak{T}_{\mathrm{PC}}(\widehat{\widetilde{\boldsymbol{F}}},\;\boldsymbol{C})\rightarrow\widehat{\boldsymbol{F}}$.
11:     Obtained target text via $\mathfrak{T}_{\mathrm{S}}^{-1}(\widehat{\boldsymbol{F}})\rightarrow\widehat{\widetilde{\boldsymbol{T}}}$.
12:     Compute ${\mathcal{L}}_{\mathrm{PN}}(\boldsymbol{\delta})$ and ${\mathcal{L}}_{\mathrm{PC}}(\boldsymbol{\zeta})$ to update $\boldsymbol{\delta}$ and $\boldsymbol{\zeta}$.
13:end while
14:Output: Trained ${\mathfrak{T}}_{\mathrm{PN}}(\cdot)$ and ${\mathfrak{T}}_{\mathrm{PC}}(\cdot)$.
According to the compensation operation at the receiver, the understandability of translated text, $\widehat{\widetilde{\boldsymbol{T}}}$ can be improved compared to scenarios where the received features are directly fed into the deep semantic decoder. The details for training the probe network and the probe-aided compensator are introduced in Algorithm 3.
As shown in Fig. 2 (c), the trained GAN generator and probe network are invoked to acquire features $\widetilde{\boldsymbol{F}}$ and semantic nose probe $\boldsymbol{C}$, respectively, according to the corrupted speech input. In addition, the trained channel encoder and channel decoder from the first training stage are utilized to enable robust semantic communications. The probe-aided compensator from the second training stage calibrates the corrupted deep semantic features, and the deep semantic decoder generates text in the target language.
@@ IV Numerical Results
In the experiments, the corpus CoVoST 2 is used as the clear speech dataset. To create the corrupted speech dataset, the clear speech is engulfed by semantic impairments. The bilingual evaluation understudy (BLEU) and semantic textual similarity (STS) [14] are employed as the efficient performance metrics to evaluate the performance of the proposed Ross-S2T over AWGN channels and Rayleigh fading channels with the accurate channel state information (CSI). Moreover, we chose English as the source language and Chinese as the target language.
@@ IV-A Simulation Settings
The DeepSC-S2T and the GAN-enabled deep semantic compensator are trained separately to cope with the dynamic speech input. In the DeepSC-S2T, the deep semantic encoder consists of five CNN modules and eight transformer modules, and the channel encoder/decoder has three dense layers with 1024 units. Six transformer modules are utilized in the deep semantic decoder. Moreover, the generator of the Ross-S2T is constructed by seven CNN modules and two dense layers followed by eight transformer modules, and the discriminator includes six transformer modules and three CNN modules followed by a dense layer and a softmax layer. The hyperparameters $\kappa=0.95$ and $\xi=10$. The batch size is 128 and the number of warm-up steps is 4000. The details of various NN settings are summarized in Table I.
TABLE I: Parameter settings of the proposed Ross-S2T.
Layer Name Parameters Activation
Ross-S2T Deep Semantic Encoder 7$\times$1D CNN modules 512 channels None
6$\times$×Transformer Modules 128 (8 heads) GELU
DeepSC-S2T Channel Encoder 2$\times$×Dense Layers 1024 units ReLU
Channel Decoder 2$\times$×Dense Layers 1024 units ReLU
Deep Semantic Decoder 6$\times$×Transformer Modules 128 (8 heads) GELU
GAN Generator 7$\times$1D CNN modules 512 channels None
8$\times$×Transformer Modules 128 (8 heads) GELU
Discriminator 5$\times$2D CNN modules 512 channels None
1$\times$×Dense Layer 1024 units Sigmoid
Probe Network 3$\times$×Dense Layer 1024 units Sigmoid
Probe-Aided Compensator 5$\times$2D CNN modules 512 channels None
The BLEU results of the Ross-S2T are shown in Fig. 3, where the ground truth results are obtained by feeding the clear speech into an S2TT pipeline constructed from the conformer and the BART, and a benchmark is provided by a conventional speech transmission system consisting of the adaptive multi-rate code, the polar code, and the 64-QAM. From the figure, the DeepSC-S2T dramatically outperforms the conventional approach in the context of clear speech. However, the BLEU scores obtained by the DeepSC-S2T are lower than the proposed Ross-S2T under the observed channel environments when contending with the corrupted speech, verifying the superiority of the dual-compensator mechanism.
Fig. 4 presents the STS comparison of various methods. From the figure, the Ross-S2T attains the STS score of around 0.6 under the AWGN channels when SNR$=$1 dB, while the STS score of the conventional system falls below 0.3. Moreover, the DeepSC-S2T manifests a significantly inferior capability in recovering the impaired semantic information within the corrupted speech compared to the Ross-S2T.
(a) AWGN channels
(b) Rayleigh channels
Figure 3: Simulation results of BLEU scores (4-grams).
@@ V Conclusions
In this paper, we study the robust semantic communications for speech transmission, named Ross-S2T, to support end-to-end speech-to-text translation (S2TT). Particularly, a deep semantic encoder is developed to learn textual semantic features related to another language from the clear speech, which enables the deep semantic exchange to achieve S2TT at the receiver. Moreover, a generative adversarial network (GAN)-enabled deep semantic compensator and a probe-aided compensator are tailored for corrupt speech scenarios by estimating the impaired semantic information and attaining as accurate deep semantic features as possible. Simulation results demonstrated the superiority of the proposed Ross-S2T to serve S2TT tasks and suppress semantic impairments.

### v3

@@ III-B Ross-S2T
As aforementioned, the deep semantic compensator is responsible for estimating the damaged semantic information in $\overline{\boldsymbol{S}}$, extracting the deep semantic features with the least dissimilarity to $\boldsymbol{F}$, and returning the semantic impairment probe matrix to record the positional information of corrupted deep semantic features. In the second training stage, we propose a dual-compensator mechanism, including the GAN-enabled deep semantic compensator at the transmitter and the probe-aided compensator at the receiver. Particularly, the trained deep semantic encoder is leveraged to obtain $\boldsymbol{F}$ as the real data for the discriminator. The generator is developed to process the corrupted speech, $\overline{\boldsymbol{S}}$, and generate the fake data $\widetilde{\boldsymbol{F}}$ to fool the discriminator by adopting the 1D CNN module followed by the latent space, $\boldsymbol{Z}$, and the transformer module. Note that the intermediate semantic representation, $\widetilde{\boldsymbol{I}}$, are attained as the output of the 1D CNN module. The discriminator distinguishes whether the input data is real or fake, incorporating the 2D CNN module before latent space $\boldsymbol{Z}$ followed by the dense and sigmoid layers. Denote the trainable NN parameters of the discriminator and the generator as ${\boldsymbol{\gamma}}_{\mathrm{D}}$ and ${\boldsymbol{\gamma}}_{\mathrm{G}}$, respectively, i.e., $\boldsymbol{\gamma}=({\boldsymbol{\gamma}}_{\mathrm{D}},\;{\boldsymbol{\gamma}}_{\mathrm{G}})$. Then, the loss function adopted for training the discrimination can be expressed as
${\mathcal{L}}_{\mathrm{D}}(\boldsymbol{S},\;\overline{\boldsymbol{S}})=\frac{1}{2}\left({\mathfrak{T}}_{\mathrm{D}}\left({\mathfrak{T}}_{\mathrm{DS}}\left(\boldsymbol{S}\right)\right)-1\right)^{2}+\frac{1}{2}\left({\mathfrak{T}}_{\mathrm{D}}({\mathfrak{T}}_{\mathrm{G}}(\overline{\boldsymbol{S}}))\right)^{2},$ (8)
where ${\mathfrak{T}}_{\mathrm{D}}(\cdot)$ is the discriminator and ${\mathfrak{T}}_{\mathrm{G}}(\cdot)$ is the generator.
The ${\boldsymbol{\gamma}}_{\mathrm{G}}$ can be updated as follows,
$\begin{split}{\mathcal{L}}_{\mathrm{G}}(\boldsymbol{S},\;\overline{\boldsymbol{S}})&=\frac{1}{2}\xi{\mathcal{L}}_{\mathrm{MSE}}+\frac{1}{2}{\mathcal{L}}_{\mathrm{ADV}}\\ &=\frac{1}{2}\xi\left(\boldsymbol{F}-\widetilde{\boldsymbol{F}}\right)^{2}+\frac{1}{2}\left({\mathfrak{T}}_{\mathrm{D}}({\mathfrak{T}}_{\mathrm{G}}(\overline{\boldsymbol{S}}))-1\right)^{2},\end{split}$ (9)
@@
The trained generator aims to create realistic data to deceive the discriminator and produce compensated deep semantic features, $\widetilde{\boldsymbol{F}}$, that closely resemble $\boldsymbol{F}$.
Furthermore, a probe network, $\mathfrak{T}_{\mathrm{PN}}(\cdot)$, is established at the transmission, taking $\widetilde{\boldsymbol{I}}$ as the input and providing the semantic impairment probe, $\boldsymbol{C}$. The loss function for training the $\mathfrak{T}_{\mathrm{PN}}(\cdot)$ is denoted as
${\mathcal{L}}_{\mathrm{PN}}(\boldsymbol{I},\;\widetilde{\boldsymbol{I}},\;\boldsymbol{C};\;\boldsymbol{\varepsilon})=\sum_{l^{{}^{\prime}}=1}^{L^{{}^{\prime}}}\left(i_{l^{{}^{\prime}}}-c_{l^{{}^{\prime}}}{\widetilde{i}}_{l^{{}^{\prime}}}\right)^{2},$ (10)
where $\boldsymbol{I}$ is the intermediate semantic representation extracted from the clear speech. $\boldsymbol{\varepsilon}$ is the NN parameters of $\mathfrak{T}_{\mathrm{PN}}(\cdot)$.
To further enhance the fidelity of the received deep semantic features, $\widehat{\widetilde{\boldsymbol{F}}}$, the learned semantic impairment probe, $\boldsymbol{C}$, is utilized to identify the corrupted deep semantic features in $\widehat{\widetilde{\boldsymbol{F}}}$ and the CNN-enabled probe-aided compensator commits to reducing semantic errors between the identified corrupted features and the corresponding accurate features. Denote the NN parameters of the probe-aided compensator is $\boldsymbol{\zeta}$ and it can be updated by
@@
where $l^{\backprime}$ indicates the position corresponding to the semantic impairment probe with the value of one. By introducing the probe-aided compensator, the understandability of the target text can be improved compared to scenarios where the received features are directly fed into the deep semantic decoder.
As shown in Fig. 3 (c), the trained GAN generator and probe network are invoked to acquire features $\widetilde{\boldsymbol{F}}$ and semantic impairment probe $\boldsymbol{C}$, respectively, under circumstances of corrupted speech input. The probe-aided compensator calibrates the corrupted deep semantic features, and the deep semantic decoder generates the text in the target language.
@@ IV Numerical Results
In the experiments, the corpus CoVoST 2 is used as the clear speech dataset. To create the corrupted speech dataset, the clear speech is engulfed by semantic impairments, including background noise and external speech interference. The semantic textual similarity (STS) [15] is adopted to evaluate the performance of the proposed Ross-S2T over wireless channels with the accurate channel state information (CSI). Moreover, we choose English as the source language and Chinese as the target language.
@@ IV-A Simulation Settings
In the DeepSC-S2T, the semantic extractor includes seven CNN modules, and the semantic converter consists of 12 transformer modules and three CNN modules. The channel encoder/decoder has two dense layers with 1024 units. Eight transformer modules are utilized in the deep semantic decoder. Moreover, the generator of the Ross-S2T is constructed by seven CNN modules, two dense layers, 12 transformer modules, and three CNN modules. The discriminator involves five CNN modules and one dense layer. The probe network includes three dense layers and the probe-aided compensator has five CNN modules. The hyperparameters $\kappa=0.95$ and $\xi=10$.
The STS results of the Ross-S2T are shown in Fig. 4, where the ground truth results are obtained by feeding the clear speech into an S2TT pipeline constructed from the conformer and the BART. A benchmark is provided by the conventional speech transmission system consisting of the adaptive multi-rate code, the polar code, and the 16-QAM. From the figure, the Ross-S2T attains the STS score of over 0.5 under the Rayleigh channels when SNR$=$1 dB, while the STS score of the conventional system falls below 0.4. Moreover, the DeepSC-S2T manifests a significantly inferior capability in recovering the impaired semantic information within the corrupted speech compared to the proposed Ross-S2T, verifying the superiority of the developed dual-compensator mechanism.
@@ V Conclusions
In this paper, we study the robust semantic communications for speech transmission, named Ross-S2T, to support end-to-end speech-to-text translation (S2TT). Particularly, a deep semantic encoder is developed to learn textual semantic features related to another language from the clear speech, which enables the deep semantic exchange to achieve S2TT at the receiver. Moreover, a generative adversarial network (GAN)-enabled deep semantic compensator and a probe-aided compensator are tailored for corrupted speech scenarios by estimating the impaired semantic information and attaining as accurate deep semantic features as possible. Simulation results demonstrated the superiority of the Ross-S2T in serving the S2TT task and suppressing the semantic impairments.